OpenStudy (anonymous):

The weight of an object varies inversely as the square of the distance from the center of the earth. At sea level (3978 mi from the center of the earth), an astronaut weighs 118 lb. Find her weight when she is 412 mi above the surface of the earth and the spacecraft is not in motion.

OpenStudy (anonymous):

W = k/r^2 k = 118*3978^2 W = (118*3978^2) / (3978+412)^2 = 96.8908 lbs

OpenStudy (dumbcow):

first set up the inverse relationship w = k/d^2, where k is a constant plug in the given info to find k 118 = k/(3978)^2 k = 118*(3978)^2 Then use it to find the new weight as done above ^^

OpenStudy (anonymous):

thank you so much